A plane took 1 hour longer to travel 560 miles on the first portion of a flight than it took to fly 480 miles on the second portion. If the speed was the same for each portion, what was the flying time for the second part of the trip?
I am so stumped on this one I have not got a clue where to begin. Please give me a hand Thank you

Answer by stanbon(48554)

1st leg Data: dist=560; time=x+1; rate=d/t=560/(x+1)
2nd leg Data: dist=480; time=x; rate=d/t=480/x
EQUATION:
rate = rate
560/(x+1)=480/x
560x=480(x+1)
560x=480x+480
80x=480
x=6 hrs (time for 2nd leg of the trip)
Cheers,
stan H.